Given that the joint probability distribution of the discrete random variables X and Y is given by



f(x,y)= { k(4x+3y) x=1,2,3 y=1,2

{ 0 otherwise

Find

i) The value of the constant k

ii) The marginal probability distribution function X

iii)The conditional probability distribution Y given x=2

iv) The conditional mean given X=2


1
Expert's answer
2021-10-13T07:52:55-0400

i)


"X""Y\ \def\arraystretch{1.5}
 \begin{array}{c:c:c:c}

 & 1 & 2 & 3 \\ \hline
 1 & 7k & 11k & 15k \\
 \hdashline
 2 & 10k & 14k & 18k
\end{array}"

"7k+11k+15k+10k+14k+18k=1"

"k=\dfrac{1}{75}"

ii)


"Y\ \def\arraystretch{1.5}
 \begin{array}{c:c:c:c}

 & 1 & 2 & 3 \\ \hline
 1 & 7/75 & 11/75 & 15/75\\
 \hdashline
 2 & 10/75 & 14/75 & 18/75
\end{array}"

"X=1:"


"P(X=1|Y=1)+P(X=1|Y=2)"

"=7\/75+10\/75=17\/75"

"X=2:"


"P(X=2|Y=1)+P(X=2|Y=2)"

"=11\/75+14\/75=25\/75=1\/3"

"X=3:"


"P(X=3|Y=1)+P(X=3|Y=2)"

"=15\/75+18\/75=33\/75=11\/25"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 x & 1 & 2 & 3 \\ \hline
 f_X(x) & 17/75 & 1/3 & 11/25 
\end{array}"

iii)


"P(Y=1|X=2)=11\/25"

"P(Y=2|X=2)=14\/25"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 y & 1 & 2 \\ \hline
 f_{Y|X=2}(y) & 11/25 & 14/25 
\end{array}"

iv)


"E(Y|X=2)=1(11\/25)+2(14\/25)=39\/25"